thirty-five students went out for the team. this was five-eights of the students in the grade.
a. how many students in the grade?
b. what percent of the class went out for the team?

Answer by rothauserc(4718)

a) let x be the number of students in the grade
:
(5/8) * x = 35
:
5x = 280
:
x = 56
:
*******************************************
there are 56 students in the grade
*******************************************
:
b) 35/56 = 7/8 of the class went out for the team
:
Note we assume grade and class are equivalent
: